Please give explanation to these

8.Which one of the following is a working capital management decision?

Answers:

Should a new machine be purchased this year?

Correct

Should the accounts payables level be decreased?

Should debt or equity financing be used to purchase a building?

Should 10 or 15-year bonds be issued?

29.


Other things held constant, for a given change in the market interest rate, the _______the time to maturity of bond, the ____ the change in the bond price.

Correct answer:

shorter; smaller.

33.

A store is offering a diamond ring for sale for 24 months at $100 per month. The retail price of the ring is $2,800. What is the interest rate on this offer?

Correct answer:

14.38%

34.

What’s the value of a 30-year, $1,000 par value, 6% coupon rate bond if the yield to maturity (YTM) is 6%?

Correct answer:

$1,000

45

A company is expected to pay their first annual dividend at the end of this year. That payment will be $0.50 a share. The company will increase the dividend by 4% per year in the future. The required return is 12%. What is the estimated value of this stock today?

Correct answer:

$6.25

49
Answer #1

1.
Working capital is related to current assets or current liabilities
and out of the given options, accounts paybales is current liabilities

2.
Interest rate sensitivity is high for high duration and duration is directly proportional to maturity
So, shorter the maturity smaller is the change in the bond price

3.
Using financial calculator
PMT=-100
N=24
PV=2800
FV=0
CPT I/Y=1.20%

Hence, interest rate=1.20%*12=14.40%

4.
As yield to maturity is equal to coupon rate, the price must be equal to par, price=1000

5.
=D1/(r-g)
=0.50/(12%-4%)
=6.25